OpenStudy (anonymous):

Solve for x: 1 - 2 square root x = 11 -4 square root x

OpenStudy (anonymous):

simply by getting the x parts together on one side and the other stuff on the other side of an equation. then square both sides. be sure to check all solutions in the original equation for any extraneous roots.

OpenStudy (anonymous):

\(1-2\sqrt{x}=11-4\sqrt{x}\) this is your equation, yeah?

OpenStudy (anonymous):

Yes that's the equation

OpenStudy (anonymous):

okay, so what's your first move?

OpenStudy (anonymous):

So you said get all the x parts together

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

So I add -2 square root x to the other side, sorry im really bad at math

OpenStudy (anonymous):

first, stay positive. if you tell yourself and others that you're bad at math, guess what? you will be! You're obviously working to improve, right? so say that! 2nd, think about eliminating the x term from one side of the eqn. left or right side, it doesn't matter. now how would you do that? if you eliminate it from the left side (as I think you are attempting to do), you must add the opposite of what's there. What is the opposite of \(-2\sqrt{x}\) ?

OpenStudy (anonymous):

Ok, thank you! The opposite would be positive

OpenStudy (anonymous):

very good!!! so what you do to one side, you have to do to the other (in order to maintain equality). make sense? so what do you get?